A gas, while expanding under isobaric conditions, does 465 J of work. The pressure of the gas is 1.90×1051.90×105 Pa, and its initial volume is 1.50×10−31.50×10−3m³. What is the final volume of the gas?
Answer #1

formula for work done is

W = P del V

W = P ( Vf- Vi)

465 = 1.9* 10^5 ( Vf - 1.5* 10^-3)

Vf = 3.94 * 10^-3 m^3